OpenStudy (anonymous):

Lillian made a down payment of $150 on some hockey equipment. The balance of the purchase price was paid for using a deferred payment plan with an 18-month introductory period. The interest rate is 26.79%. If the purchase price of the hockey equipment was $975 and minimum payments of $12 were made during the deferment period, what is the balance after the deferment period? $1,011.60 $1,084.80 $1,215.60 $1,227.60

OpenStudy (anonymous):

I don't want the answer. I would really like to know how to do it though!

OpenStudy (anonymous):

amount financed 975-150 = 825

OpenStudy (anonymous):

he made 12 payments of 12 = 144

OpenStudy (anonymous):

add 6 more payments to that 12 * 18 = 216

OpenStudy (anonymous):

financed * interest - payments made = should be balance

OpenStudy (anonymous):

that didn't help I got 5.0175. thank you though

OpenStudy (anonymous):

"The balance of the purchase price" do you know what that means??

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

me neither.....

OpenStudy (anonymous):

fuhhh...

OpenStudy (anonymous):

A city’s population is declining continuously at a rate of 6.1% each year. If the population of the city is currently 435,410, what will the population of the city be in 8 years? 222,930 263,162 266,945 267,278 how bout this

OpenStudy (anonymous):

the balance of the purchase price is the amount AFTER he paid the down payment

OpenStudy (anonymous):

ok do you know how the 26% interest rate works - is it annualy? or a one time interest on the balance of purchase

OpenStudy (anonymous):

@itskaymarie02 it is a loss of 6.1% per year this also means that 93.9% of the city stay. ok so far?

OpenStudy (anonymous):

seems like an annual amount

OpenStudy (anonymous):

whered 93.9% come in?

OpenStudy (anonymous):

because total city 100% - 6.1% loss = 93.9% after one year, there was a loss of 6.1% and 93.9% were still there

OpenStudy (anonymous):

good so far...

OpenStudy (anonymous):

@JuanitaM I also think so next question would be, does the interest already count during the deferment period......

OpenStudy (anonymous):

@itskaymarie02 next, we need to do this sort of problem: 93.9% of "93.9%", how much is it? because in the second year, we need 93.9% of "93.9%" remaining people...

OpenStudy (anonymous):

that confused me.... im sorry..

OpenStudy (anonymous):

it is confusing

OpenStudy (anonymous):

I can take a simpler example: 100% , take 1/4 how much is this?

OpenStudy (anonymous):

25% ?

OpenStudy (anonymous):

yes now, again take 1/4 of what is still left

OpenStudy (anonymous):

of the 93.9%?

OpenStudy (anonymous):

of the 25%

OpenStudy (anonymous):

6.25?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now what this means is that: 1/4 of 1/4 of 100 =6.25%

OpenStudy (anonymous):

"1/4 of 100" =25% 1/4 of this remaining "1/4 of 100" =6.25%

OpenStudy (anonymous):

ok good so far..

OpenStudy (anonymous):

ok, so we can observe that instead of taking 1/4 of 1/4 we can also take 1/4 x 1/4 of 100 what's 1/4 x 1/4 ?

OpenStudy (anonymous):

1/16 ?

OpenStudy (anonymous):

yes :) 

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

than what... on the problem... like how do I figure it out?

OpenStudy (anonymous):

so when we have a problem like take 1/4 of 1/4 of 100 we can also just multiply everything together!

OpenStudy (anonymous):

the problem states that 6.1% of the population leaves, and they ask what population the city still has. because they ask for the city population we don't really care about the slices being carried away: we care what is still left

OpenStudy (anonymous):

the total city is 100% and 6.1% are leaving so 100%-6.1% are still there

OpenStudy (anonymous):

93.9% are still there

OpenStudy (anonymous):

ok im good on that

OpenStudy (anonymous):

now the rest :)

OpenStudy (anonymous):

the city did not stop losing population after one year though? they CONTINUE to lose 6.1% of what is left

OpenStudy (anonymous):

so they CONTINUE to retain 93.9% of whatever they still have after 1 y

OpenStudy (anonymous):

if they lose another 6.1% than they'll have 87.8 % ?

OpenStudy (anonymous):

if they lost another 6.1% of the original population, then yes HOWEVER they are going to lose 6.1% of the REMAINING POPULATION

OpenStudy (anonymous):

basically how much 6.1% is changes each year because the total population decreases each year and so the people for 6.1% are adjusted I will just use the numbers now:

OpenStudy (anonymous):

lol thank you!(:

OpenStudy (anonymous):

city had 435,410 after 1 year, they have 93.9% of 435,410 = 408,849

OpenStudy (anonymous):

then, when year 2 just begins, they adjust what 6.1% is. previously, 6.1% of 435,410 was 26,560 people. Now, 6.1% of 408,849 are 24,939 people Basically, if the same percentage of people want to leave, they're less in total numbers because there aren't that many people who're even there!

OpenStudy (anonymous):

alrighttt.. so how do I figure out the answerrr

OpenStudy (anonymous):

if 93.9% stay in year 2, that's less people than when 93.9% stay in year 1

OpenStudy (anonymous):

no we have to calculate how many stay :)

OpenStudy (anonymous):

yes 26560 are leaving so 408,849 are still there

OpenStudy (anonymous):

I wanted to calculate with the 6.1% leaving people as basis, this is not possible. because they can move anywhere they are no longer a number we can use

OpenStudy (anonymous):

6.1% x 408849 = 24,939

OpenStudy (anonymous):

24,939 people leave in the second year

OpenStudy (anonymous):

that means 408849 - 24939 people stay = 383,910 stay

OpenStudy (anonymous):

383,910 people still in city after 2 years

OpenStudy (anonymous):

the problem asks us: how many people are 93.9% of 93.9% of 93.9% ...of 435,410

OpenStudy (anonymous):

so 93.9% x 93.9% x 93.9% ?